Antoine needs to buy metal primer to rust proof a cylindrical fuel storage tank. The tank has a diameter of 15 ′ 6 ″ and a height of 18 ′ . A 3.8 L can of metal primer costs $47.13 and covers 40 m 2 . What will be the total cost of the metal primer? (Assume that the bottom of the storage tank will be painted.)
1
Expert's answer
2013-05-22T10:05:48-0400
First we find the area of ​​a cylinder:

S = 2 π rh
S = 2 * 3.14 * (15.6 / 2) * 18 = 881.712 m2

Then we find the number of cells with a capacity of 3.8 liters required:
n = 881.712 m2 / 40 m2 = 22 banks.
Cost = $ 47.13 * 22 = $ 1039

Answer: $ 1039